OpenStudy (awhite):

I need help with answering this please. I do not want the direct question, but how to obtain the correct answer. Thank you in advance.

OpenStudy (awhite):

Kyra is using rectangular tiles of two types for a floor design. A tile of each type is shown below: Two rectangular tiles, rectangle PQRS with vertices at P(2, 2), Q(4, 2), R(4, 8), and S(2, 8). Rectangle JKLM with vertices J(5, 2) , K(6, 2), L(6, 5) and M(5, 5) Which statement is correct? A). The two tiles are not similar because segment SP is to segment SR is 5 : 1 and segment MJ is to segment ML is 2 : 3. B.) The two tiles are similar because segment PQ is to segment QR is 1 : 3 and segment JK is to segment KL is also 1 : 3. C.) The two tiles are similar because segment SR is to segment SP is 3 : 2 and segment ML is to segment MJ is also 3 : 2. D.)The two tiles are not similar because segment PQ is to segment QR is 1 : 6 and segment JK is to segment KL is 1 : 3.

OpenStudy (mathstudent55):

To know if two polygons are similar, you must do two things. 1. Each angle of the first polygon must be congruent to each corresponding angle of the second polygon. 2. All the ratios of the lengths each pair of corresponding sides are equal.

OpenStudy (mathstudent55):

Do you understand the two conditions above?

OpenStudy (awhite):

Kind of.

OpenStudy (mathstudent55):

Here is an example. 

OpenStudy (mathstudent55):

The example above uses triangles, but it's the same idea. The question is: Are triangles ABC and DEF similar?

OpenStudy (mathstudent55):

To answer we need to look at the angles and at the ratios of the lengths of the sides.

OpenStudy (awhite):

Yes, because they have the same shape and not the same size...Right?

OpenStudy (mathstudent55):

Angles A and D are congruent. Angles B and E are congruent. Angles C and F are congruent. So far the first condition is met. The three angles of one triangle are congruent to the three corresponding angles of the other triangle.

OpenStudy (awhite):

Okay.

OpenStudy (mathstudent55):

Now let's look at the lengths of corresponding sides. Here are the three pairs of corresponding sides. AB and DE BC and EF AC and DF Now we need to see if the ratios of lengths are the same.

OpenStudy (awhite):

That is what I am confused about...the ratio part of the question.

OpenStudy (mathstudent55):

Ratios of lengths of corresponding sides: AB to DE = \(\dfrac{1}{3}\) BC to EF = \(\dfrac{\sqrt 3}{3 \sqrt 3} = \dfrac{1}{3}\) AC to DF = \(\dfrac{2}{6} = \dfrac{1}{3} \)

OpenStudy (mathstudent55):

You see how every side of triangle ABC is 1/3 the length of the corresponding side of triangle DEF?

OpenStudy (mathstudent55):

That is what equal ratios of the lengths of corresponding sides means.

OpenStudy (awhite):

Yes. But how do I get that with my question? How do I solve to find the ratio with coordinates?

OpenStudy (mathstudent55):

Now we need to apply this same method to your two polygons.

OpenStudy (mathstudent55):

Remember the 2 steps for similarity of polygons. 1. congruent corresponding angles 2. equal ratios of corresponding side lengths

OpenStudy (mathstudent55):

Let's look at the angles of the two polygons. Can you tell what angles these two quadrilaterals have?

OpenStudy (mathstudent55):

For example, look at points P and S. Do you see that the both have x-coordinate 2? What does that tell you about segment PS?

OpenStudy (awhite):

The ratio is 2? (I am really bad at this, I apologize)

OpenStudy (mathstudent55):